OpenStudy (anonymous):

(9a^6 b^6) over (6a^6 b^10) divided by (3a^4 b^12) over (a^2 b^4)

OpenStudy (whpalmer4):

Ah, now I see how to do multi-digit exponents! Let's try that again. \[(9a^6b^6/6a^6b^{10})/(3a^4b^{12}/a^2b^4)\] Simplify the numerator and denominator by crossing out common factors, giving \[(9/6b^4)/(3a^2b^8)=9/(6b^4*3a^2b^8)\] which simplifies to \[9/(18a^2b^{12}) = 1/(2a^2b^{12})\] or \[a^{-2}b^{-12}/2\]